x - 2y = -1
-x + 5y = 9

Answer by mananth(16949)

x - 2y = -1
-x + 5y = 9
..
add the two equations to eliminate x
x-2y-x+5y=-1+9
3y=8
divide by 3
3y/3 = 8/3
y= 8/3
..
plug value of y in any equation
x-2y=-1
x-2*(8/3)=-1
x-16/3 =-1
x= 16/3 -1
x= (16-3)/3
x=13/3
..
-x+5y=9
(-13/3)+5*(8/3)= 9